# Towards Understanding Variants of Invariant Risk Minimization through the Lens of Calibration

### v2

@@ 1 Introduction
In the evolving landscape of machine learning, the importance of out-of-distribution (OOD) generalization (Vapnik, 1991) and calibration (Guo et al., 2017a) cannot be understated, especially in real-world applications and critical scenarios. Traditional machine learning approaches, grounded in the assumption that data are independently and identically distributed (IID), often struggle to cope with OOD scenarios. Additionally, a notable trend in contemporary machine learning models is their tendency towards overconfidence, which undermines the reliability of their confidence estimations. This issue is primarily recognized as a calibration challenge. Responding to these limitations, there has been a surge in research focusing on methodologies like Invariant Risk Minimization (IRM) (Arjovsky et al., 2020). IRM is designed to identify constant features across different environments, facilitating more robust generalization in the face of environmental variations. However, the computational demands of IRM have led to the exploration of more feasible, approximate methods. Despite the development of several IRM variants, their inability to consistently surpass the performance of finely-tuned Empirical Risk Minimization (ERM) models has been observed (Gulrajani & Lopez-Paz, 2020)(Zhang et al., 2023a). This performance gap raises critical questions about the inherent deficiencies of these IRM variants. As we showed in Figure 1, even by leveraging IRM variants, there is a trade-off between in-distribution (ID) and OOD accuracy. It implies that models trained by IRM variants failed to obtain invariant features since they sacrifice ID accuracy to achieve better OOD accuracy. From another perspective of IRM, it can also be conceptualized as a specialized form of multi-domain calibration (Wald et al., 2021a). Our study intersects with these findings, examining how IRM variants mimic the original model in calibration. We posit that effective implementation of IRM should lead to reduced loss and a zero expected calibration error (ECE) (Wald et al., 2021a). Through comparative experimentation, we assess various IRM variants in environments marked by distributional shifts, such as diversity and correlation shifts (Ye et al., 2021). Our evaluation focuses not only on accuracy but also on ECE. The results highlight notable disparities in ECE among different methodologies, even when OOD accuracy levels are similar. This suggests that a sole reliance on OOD accuracy for method development might be inadequate. Our observations include a range of calibration behaviors across methods, from overconfidence to underestimation.
@@
  • We used ECE to analyze discrepancies between IRM formulation and its variants.
  • We observed that Information Bottleneck-based IRM (IB-IRM) effectively lowers ECE, aligning it more closely with the original objectives of IRM (Figure 3, Table 2).
  • We reveal that IRM variants with robust regularization typically achieve lower ECE (Figure 3).
  • We identify a trade-off between ECE and accuracy, thereby underscoring the intricate balance and compromises necessary for achieving both robust generalization and accurate calibration (Figure 4).
@@ 2.1 OOD Generalization
However, in the real world, the assumption of IID does not always hold, and distribution shifts between training and testing environments can occur(Quinonero-Candela et al., 2008). When the model relies on features that are only effective in the training environment, known as shortcut features, its performance can decrease in the testing environment(Geirhos et al., 2020). This issue is commonly known as OOD generalization and has become an urgent problem recently. Therefore, it is necessary to train robust models that are not dependent on specific environments and can perform consistently across all environments. A robust model is defined as follows.
@@
where $E_{all}$ represents the set of all possible environments. (2) indicates that, for a robust model, the conditional probability of the label given the model’s prediction should be equal between any two environments.
Since it is difficult to access all real-world environments during training, various methods have been proposed to realize invariant models satisfying (2) using limited learning environments.
@@ 2.2 Calibration
Equation (3) implies that the model’s predictive probability is always consistent with its actual accuracy.
@@ 2.3 OOD Generalization and Calibration
It has been pointed out that OOD generalization and calibration are interrelated. Notably, (3) can be a sufficient condition for (2), theoretically demonstrating that calibration across all environments leads to OOD generalization. Wald et al. (2021b) capitalize on this observation and propose CLOvE, a method that calibrates models across multiple environments, thereby enabling generalization to unseen data distributions.
Additionally, it has been empirically demonstrated that there is a correlation between a model’s accuracy in OOD scenarios and its calibration performance. It has been demonstrated that there exists a correlation between a model’s accuracy and its ECE under dataset shift scenarios. Furthermore, techniques such as temperature scaling(Guo et al., 2017b), originally proposed for calibration purposes, have been shown to be effective for OOD generalization.(Ovadia et al., 2019; Naganuma & Hataya, 2023).
In this paper, we utilize this fact and hypothesize that models with high calibration performance will likely achieve better OOD generalization. We evaluate the approximation methods of Invariant Risk Minimization (IRM), introduced in 4, based on the model’s calibration performance to assess their OOD generalization capabilities.
@@ 3.1 Domain Invariant Representation
In the context of OOD scenarios, there is an application for Representation Learning, specifically aimed at extracting environment-invariant features from raw data and reducing environment-specific spurious features to achieve more robust predictions. For example, in a task of classifying animals from images, if the model learns to associate parts of an animal with extraneous elements, it may lead to high performance within the training domain but fail to make robust predictions when the background distribution shifts in the test domain. However, if the model discards spurious features (e.g., background) from the raw data, it can consistently make robust predictions across all environments. One particularly well-known method that applies representation learning to OOD generalization scenarios is Domain-Adversarial Neural Networks (DANN) (Ganin et al., 2016). This approach uses adversarial training to adjust the learning process so that it is impossible to discern from the extracted features which environment the input comes from, thereby aiming to extract invariant features.
@@ 3.2 Invariant Learning
In the context of applying representation learning to OOD generalization, the method we focus on in this paper is IRM(Arjovsky et al., 2020). IRM is proposed with the aim of extracting environment-invariant features from input data to enable consistent predictions across any environment. IRM is defined with the purpose of realizing (2) as follows.
@@
Here, $\mathcal{X}$ represents the input space of each data point, $\widehat{\mathcal{H}}$ denotes the extracted invariant feature space, and $\mathcal{Y}$ is the output space of the model. IRM defines the function as $f=\omega\circ\Phi$, where $\Phi$ maps the input data to $\widehat{\mathcal{H}}$, and based on these extracted invariant features, the final prediction is made by $\omega$ The aim of IRM is to achieve consistent predictions as a form of OOD generalization.
The formulation in (5) represents a bi-level optimization problem, which is challenging to solve exactly. As a result, various implementable approximation methods have been proposed. Many variants add various constraints to the IRM objective and approximate the bi-level optimization problem as a single-level problem to make it more tractable (Arjovsky et al., 2020; Ahuja et al., 2022; Chen et al., 2022; Ahuja et al., 2020; Lin et al., 2022). Additionally, Zhang et al. (2023b) propose an improved variant of IRM Game, another IRM variant, that retains the bi-level optimization framework during training. However, despite numerous contributions to the development of approximation methods for IRM, a method that generally surpasses well-tuned ERM in terms of OOD generalization accuracy has not yet been proposed. Therefore, it is essential to understand why these approximation methods have not succeeded, and what approaches might be more suitable for addressing the approximation problem.
While the original IRM formulation poses implementation challenges, a alternative method has also been proposed to address this issue. IRM aims to learn invariant features by keeping $\mathbb{E}[Y^{e}|f(X^{e})]$ constant across environments. However, Huh & Baidya (2023) introduced a complementary notion of invariance called MRI, which seeks to conserve the label-conditioned feature expectation $\mathbb{E}[f(X^{e})|Y^{e}]$ across environments, in contrast to the original IRM formulation.
@@ 4.1 IRMv1
IRMv1(Arjovsky et al., 2020) constrains $\omega$ in (5) as a linear mapping and allows the transformation $\Phi^{\prime}=(\omega\cdot\Phi)$, $\omega^{\prime}=1.0$. The following one-variable optimization problem can be relaxed and computed.
@@ 4.4 IRM Game
IRM Game(Ahuja et al., 2020) aims to achieve Nash equilibrium in terms of inference accuracy by introducing a game-theoretic framework to the IRM between each environment. Each environment $e$ in the training data is assigned its own predictor $\omega^{e}$, each of which is trained to be optimal in each environment, and the final $\omega_{ens}$ is the ensemble of all predictors $\frac{1}{|E_{train}|}\sum_{e\in E_{train}}\omega^{e}$. By learning environment-specific predictors, the limitation on linearity made in IRMv1 is eliminated, and an implementation closer to (5) is aimed at.
@@ 5.1 Setting
For optimization, Adam was used consistently across all models, and the tuning of learning rates and hyperparameters for each method was conducted in accordance with their respective papers.
@@ 5.2.2 Results
Figure 3: Comparison of the model’s OOD ECE in CMNIST, with various values of $\lambda$ in the formulation of IRMv1 6. It was observed that increasing $\lambda$, thereby increasing the regularization penalty in IRMv1, results in a lower ECE for the model in the OOD environment.
Fig. 2 presents a scatter plot showing the relationship between OOD accuracy and ECE for multiple IRM variants, each trained with multiple hyperparameters. The red solid line represents the theoretical value when all model’s predictive probabilities are 100%, indicating very high confidence, while the blue solid line represents the case when all predictive probabilities are 50%, indicating very low and ambiguous confidence. The methods with relatively high accuracy (IRMv1(green), IB-IRM(light blue), BIRM(red)) align with the blue solid line, showing that they can overcome excessive fitting to non-invariant features by making predictions more ambiguous. On the other hand, ERM(blue) and IRM Game(purple) tend to be distributed near the red solid line, indicating predictions with relatively high confidence. PAIR(yellow) is distributed between the two solid lines, demonstrating characteristics that are intermediate between the two aforementioned groups.
Figure 3 visualizes how the penalty introduced in IRMv1 for the purpose of OOD generalization actually affects the ECE. It specifically compares different values of $\lambda$ in IRMv1’s formulation 6, demonstrating that as the impact of IRMv1’s penalty increases, the ECE decreases, suggesting better calibration.
@@ 5.3.2 Results
The average results from three different random seeds under the aforementioned conditions are presented in Table 2, where although all methods achieved a similar average OOD accuracy, notable differences were observed in OOD ECE. Specifically, the ECE of IB-IRM was, on average, lower than that of any other method. Additionally, Figure 5(c) visualizes the relationship between ECE in the training and test environments, with the red solid line representing cases where ECE is equal in both environments. Ideally, models should be distributed near this red line from the perspective of cross-environmental calibration of uncertainty. Specifically, in Figures 5(a) and 5(c), IB-IRM is distributed near the red solid line compared to other methods, suggesting that it does not overfit to the training environment and exhibits more ideal feature learning.
@@ 6 Information Bottleneck in Calibration
First, we investigated the impact of the penalty term related to the information bottleneck in IBIRM on calibration. Figure 6 visualizes various evaluation metrics in CMNIST when altering the strength of the information bottleneck in IBIRM, that is, the magnitude of $\gamma$ in (7). As shown in panel (a), which visualizes this based on OOD ECE and OOD Accuracy, we confirm the clear trend that strengthening the regularization of $\gamma$ improves both accuracy and ECE. Panel (b) shows the relationship between ECE in ID and OOD. Increasing $\gamma$ aligns both ECEs towards the red line, representing equality for conditional probability. In other words, this constitutes a successful calibration across environments. These findings suggest that not only is the IRMv1’s regularization term aimed at OOD scenarios, but the information bottleneck also further promotes the learning of invariant features.
Additionally, we demonstrate the validity of our results based on the theoretical principles of the information bottleneck. In OOD scenarios, the objective of the information bottleneck is to minimize the mutual information $I(X;\Phi(X))$ between the input $X$ and the intermediate representation $\Phi(X)$ while preserving the mutual information $I(\Phi(X);Y)$ between $\Phi(X)$ and the output $Y$. Intuitively, this means learning $\Phi(X)$ to reduce as much information from $X$ as possible while retaining as much information about $Y$. When $\Phi(X)$ is a deterministic transformation of $X$, the entropy $H(\Phi(X))$ can be used to minimize $I(X;\Phi(X))$ (Kirsch et al., 2021). In the context of IRM, which aims to learn invariant features, keeping $H(\Phi(X))$ small encourages $\Phi$ to capture invariant features $X_{inv}$ from $X$ and discard spurious features $X_{sup}$. In fact, Ahuja et al. (2022) theoretically demonstrated, using realistic SEMs, that minimizing $H(\Phi(X))$ in IRM leads to $\Phi$ discarding spurious features and grasping invariant features. Our results empirically support the effectiveness of applying the information bottleneck method to invariant feature learning.
@@
Figure 6: The figures illustrate the impact of the information bottleneck on calibration within IBIRM. Panel (a) displays the relationship between ECE and Accuracy in the OOD context, investigated by varying the coefficient $\gamma$ of the information bottleneck penalty in the IBIRM formulation (7). As the value of $\gamma$ increases, both metrics show improvement. Panel (b) visualizes the values of ECE in both ID and OOD. Same as (a), altering the value of $\gamma$ shows that the larger the value, the more the data aligns with the red line, which indicates equality between the two ECEs. With a sufficiently large $\gamma$, the points are almost perfectly distributed along the red line, indicating successful calibration across multiple environments.

### v3

@@ 1 Introduction
In the evolving landscape of machine learning, the importance of out-of-distribution (OOD) generalization (Vapnik, 1991) and calibration (Guo et al., 2017a) cannot be understated, especially in real-world applications and critical scenarios. Traditional machine learning approaches, grounded in the assumption that data are independently and identically distributed (IID), often struggle to cope with OOD scenarios. Additionally, a notable trend in contemporary machine learning models is their tendency towards overconfidence, which undermines the reliability of their confidence estimations. This issue is primarily recognized as a calibration challenge. Responding to these limitations, there has been a surge in research focusing on methodologies like Invariant Risk Minimization (IRM) (Arjovsky et al., 2020). IRM is designed to identify constant features across different environments, facilitating more robust generalization in the face of environmental variations. However, the computational demands of IRM have led to the exploration of more feasible, approximate methods. Despite the development of several IRM variants, their inability to consistently surpass the performance of finely-tuned Empirical Risk Minimization (ERM) models has been observed (Gulrajani & Lopez-Paz, 2020; Zhang et al., 2023a). This performance gap raises critical questions about the inherent deficiencies of these IRM variants. As we showed in Figure 1, even by leveraging IRM variants, there is a trade-off between in-distribution (ID) and OOD accuracy. This implies that models trained by IRM variants failed to obtain invariant features since they sacrificed ID accuracy to achieve better OOD accuracy. From another perspective of IRM, it can also be conceptualized as a specialized form of multi-domain calibration (Wald et al., 2021a). Our study intersects with these findings, examining how IRM variants mimic the original model in calibration. We posit that effective implementation of IRM should lead to reduced loss and a zero expected calibration error (ECE) (Wald et al., 2021a). Through comparative experimentation, we assess various IRM variants in environments marked by distributional shifts, such as diversity and correlation shifts (Ye et al., 2022). Our evaluation focuses not only on accuracy but also on ECE. The results highlight notable disparities in ECE among different methodologies, even when OOD accuracy levels are similar. This suggests that a sole reliance on OOD accuracy for method development might be inadequate. Our observations include a range of calibration behaviors across methods, from overconfidence to underestimation.
@@
  • We study discrepancies between IRM formulation and its implemented variants using ECE as a key metric to measure how much the model learned environmental invariant features.
  • We reveal that IRM variants with strong regularization typically achieve lower ECE (Figure 3 and Figure 6).
  • We observed that Information Bottleneck-based IRM (IB-IRM) lowers ECE across the environments (Figure 5(c)), aligning it more closely with the original objectives of IRM while maintaining OOD accuracy relatively (Figure 2 and Table 2).
  • In contrast, for IRMv1, we identify a trade-off between ECE and accuracy, thereby underscoring the compromises necessary for achieving robust generalization and accurate calibration (Figure 4).
@@ 2.1 OOD Generalization
However, in the real world, the assumption of IID does not always hold, and distribution shifts between training and testing environments can occur (Quinonero-Candela et al., 2008). When the model relies on features that are only effective in the training environment, known as shortcut features, its performance can decrease in the testing environment (Geirhos et al., 2020). This issue is commonly known as OOD generalization and has become an urgent problem recently. Therefore, it is necessary to train robust models that are not dependent on specific environments and can perform consistently across all environments. A model that is robust across the environment is defined as follows.
@@
where $E_{all}$ represents the set of all possible environments. Equation 2 indicates that, for a robust model, the conditional probability of the label given the model’s prediction should be equal between any two environments.
Since it is difficult to access all real-world environments during training, various methods have been proposed to realize invariant models satisfying (eq. 2) using limited learning environments.
@@ 2.2 Calibration
Equation 3 implies that the model’s conditional predictive probability (in other words, it is called confidence) is always consistent with its actual accuracy.
@@ 2.3 OOD Generalization and Calibration
It has been pointed out that OOD generalization and calibration are interrelated. Notably, eq. 3 can be a sufficient condition for eq. 2, theoretically demonstrating that calibration across all environments leads to OOD generalization. Wald et al. (2021b) capitalize on this observation and propose CLOvE, a method that calibrates models across multiple environments, thereby enabling generalization to unseen data distributions.
Additionally, it has been empirically demonstrated that there is a correlation between a model’s accuracy in OOD scenarios and its calibration performance. It has been demonstrated that there exists a correlation between a model’s accuracy and its ECE under dataset shift scenarios. Furthermore, techniques such as temperature scaling (Guo et al., 2017b), originally proposed for calibration purposes, have been shown to be effective for OOD generalization (Ovadia et al., 2019; Naganuma & Hataya, 2023).
In this paper, we utilize this fact and hypothesize that models with high calibration performance will likely achieve better OOD generalization. We evaluate the approximation methods of Invariant Risk Minimization (IRM), introduced in Section 4, based on the model’s calibration performance to assess their OOD generalization capabilities.
@@ 3.1 Domain Invariant Representation
In the context of OOD scenarios, there is an application for representation learning, specifically aimed at extracting environment-invariant features from raw data and reducing environment-specific spurious features to achieve more robust predictions. For example, in the task of classifying animals from images, if the model learns to associate parts of an animal with extraneous elements, it may lead to high performance within the training domain but fail to make robust predictions when the background distribution shifts in the test domain. However, if the model discards spurious features (e.g., background) from the raw data, it can consistently make robust predictions across all environments. One particularly well-known method that applies representation learning to OOD generalization scenarios is Domain-Adversarial Neural Networks (DANN) (Ganin et al., 2016). This approach uses adversarial training to adjust the learning process so that it is impossible to discern from the extracted features which environment the input comes from, thereby aiming to extract invariant features.
@@ 3.2 Invariant Learning
In the context of applying representation learning to OOD generalization, the method we focus on in this paper is IRM (Arjovsky et al., 2020). IRM is proposed with the aim of extracting environment-invariant features from input data to enable consistent predictions across any environment. IRM is defined with the purpose of realizing (eq. 2) as follows.
@@
Here, $\mathcal{X}$ represents the input space of each data point, $\widehat{\mathcal{H}}$ denotes the extracted invariant feature space, and $\mathcal{Y}$ is the output space of the model. IRM defines the function as $f=\omega\circ\Phi$, where $\Phi$ maps the input data to $\widehat{\mathcal{H}}$, and based on these extracted invariant features, the final prediction is made by $\omega$ IRM aims to achieve consistent predictions as a form of OOD generalization.
The formulation in eq. 5 represents a bi-level optimization problem, which is challenging to solve exactly. As a result, various implementable approximation methods have been proposed. Many variants add various constraints to the IRM objective and approximate the bi-level optimization problem as a single-level problem to make it more tractable (Arjovsky et al., 2020; Ahuja et al., 2022; Chen et al., 2022; Ahuja et al., 2020; Lin et al., 2022). Additionally, Zhang et al. (2023b) propose an improved variant of IRM Game, another IRM variant, that retains the bi-level optimization framework during training. However, despite numerous contributions to the development of approximation methods for IRM, a method that generally surpasses well-tuned ERM in terms of OOD generalization accuracy has not yet been proposed. Therefore, it is essential to understand why these approximation methods have not succeeded and what approaches might be more suitable for addressing the approximation problem.
While the original IRM formulation poses implementation challenges, an alternative method has also been proposed to address this issue. IRM aims to learn invariant features by keeping $\mathbb{E}[Y^{e}|f(X^{e})]$ constant across environments. However, Huh & Baidya (2023) introduced a complementary notion of invariance called MRI, which seeks to conserve the label-conditioned feature expectation $\mathbb{E}[f(X^{e})|Y^{e}]$ across environments, in contrast to the original IRM formulation.
@@ 4.1 IRMv1
IRMv1 (Arjovsky et al., 2020) constrains $\omega$ in eq. 5 as a linear mapping and allows the transformation $\Phi^{\prime}=(\omega\cdot\Phi)$, $\omega^{\prime}=1.0$. The following one-variable optimization problem can be relaxed and computed.
@@ 4.4 IRM Game
IRM Game (Ahuja et al., 2020) aims to achieve Nash equilibrium regarding inference accuracy by introducing a game-theoretic framework to the IRM between each environment. Each environment $e$ in the training data is assigned its own predictor $\omega^{e}$, each of which is trained to be optimal in each environment, and the final $\omega_{ens}$ is the ensemble of all predictors $\frac{1}{|E_{train}|}\sum_{e\in E_{train}}\omega^{e}$. By learning environment-specific predictors, the limitation on linearity made in IRMv1 is eliminated, and implementation closer to eq. 5 is aimed at.
@@ 5.1 Setting
For optimization, Adam (Kingma & Ba, 2015) was used consistently across all models, and the tuning of learning rates and hyperparameters for each method was conducted in accordance with their respective papers.
@@ 5.2.2 Results
Figure 3: Comparison of the model’s OOD ECE in CMNIST, with various values of $\lambda$ in the formulation of IRMv1 eq. 6. It was observed that increasing $\lambda$, thereby increasing the regularization penalty in IRMv1, results in a lower ECE for the model in the OOD environment.
Figure 2 presents a scatter plot showing the relationship between OOD accuracy and ECE for multiple IRM variants, each trained with multiple hyperparameters. The red solid line represents the theoretical value when all model’s predictive probabilities are 100%, indicating very high confidence. In contrast, the blue solid line represents the case when all predictive probabilities are 50%, indicating very low and ambiguous confidence. The methods with relatively high accuracy (IRMv1(green), IB-IRM(light blue), BIRM(red)) align with the blue solid line, showing that they can overcome excessive fitting to non-invariant features by making predictions more ambiguous. On the other hand, ERM(blue) and IRM Game(purple) tend to be distributed near the red solid line, indicating predictions with relatively high confidence. PAIR(yellow) is distributed between the two solid lines, demonstrating intermediate characteristics between the two aforementioned groups.
Figure 3 visualizes how the penalty introduced in IRMv1 for the purpose of OOD generalization actually affects the ECE. It specifically compares different values of $\lambda$ in IRMv1’s formulation eq. 6, demonstrating that as the impact of IRMv1’s penalty increases, the ECE decreases, suggesting better calibration.
@@ 5.3.2 Results
The average results from three different random seeds under the aforementioned conditions are presented in Table 2, where although all methods achieved a similar average OOD accuracy, notable differences were observed in OOD ECE. Specifically, the ECE of IB-IRM was, on average, lower than that of any other method. Additionally, Figure 5(c) visualizes the relationship between ECE in the training and test environments, with the red solid line representing cases where ECE is equal in both environments. Ideally, models should be distributed near this red line from the perspective of cross-environmental calibration of uncertainty. Specifically, in Figure 5(a) and Figure 5(c), IB-IRM is distributed near the red solid line compared to other methods, suggesting that it does not overfit to the training environment and exhibits more ideal feature learning.
@@ 6 Information Bottleneck in Calibration
First, we investigated the impact of the penalty term related to the information bottleneck in IBIRM on calibration. Figure 6 visualizes various evaluation metrics in CMNIST when altering the strength of the information bottleneck in IBIRM, that is, the magnitude of $\gamma$ in eq. 7. As shown in panel (a), which visualizes this based on OOD ECE and OOD Accuracy, we confirm the clear trend that strengthening the regularization of $\gamma$ improves both accuracy and ECE. Panel (b) shows the relationship between ECE in ID and OOD. Increasing $\gamma$ aligns both ECEs towards the red line, representing equality for conditional probability. In other words, this constitutes a successful calibration across environments. These findings suggest that not only is the IRMv1’s regularization term aimed at OOD scenarios, but the information bottleneck also further promotes the learning of invariant features.
Additionally, we demonstrate the validity of our results based on the theoretical principles of the information bottleneck. In OOD scenarios, the objective of the information bottleneck is to minimize the mutual information $I(X;\Phi(X))$ between the input $X$ and the intermediate representation $\Phi(X)$ while preserving the mutual information $I(\Phi(X);Y)$ between $\Phi(X)$ and the output $Y$. Intuitively, this means learning $\Phi(X)$ to reduce as much information from $X$ as possible while retaining as much information about $Y$. When $\Phi(X)$ is a deterministic transformation of $X$, the entropy $H(\Phi(X))$ can be used to minimize $I(X;\Phi(X))$ (Kirsch et al., 2021). In the context of IRM, which aims to learn invariant features, keeping $H(\Phi(X))$ small encourages $\Phi$ to capture invariant features $X_{inv}$ from $X$ and discard spurious features $X_{sup}$. It should be noted that Ahuja et al. (2022) theoretically demonstrated, using realistic SEMs, that minimizing $H(\Phi(X))$ in IRM leads to $\Phi$ discarding spurious features and grasping invariant features. Our results empirically support the effectiveness of applying the information bottleneck method to invariant feature learning.
@@
Figure 6: The figures illustrate the impact of the information bottleneck on calibration with IBIRM on CMNIST task. Panel (a) displays the relationship between ECE and Accuracy in the OOD context, investigated by varying the coefficient $\gamma$ of the information bottleneck penalty in the IBIRM formulation (eq. 7). As the value of $\gamma$ increases, both metrics show improvement. Panel (b) visualizes the values of ECE in both ID and OOD. Same as (a), altering the value of $\gamma$ shows that the larger the value, the more the data aligns with the red line, which indicates equality between the two ECEs. With a sufficiently large $\gamma$, the points are almost perfectly distributed along the red line, indicating successful calibration across multiple environments.